OpenStudy (anonymous):

A small school has 100 students who occupy three classrooms. A, B and C after the first period of the school day half of the students in room A move to room B one fifth of the students in room B move to room C and one thrid of the students in room C move to room A. The total number of students in each room is the same for both periods. How many students are in each room

OpenStudy (anonymous):

solve a+b+c=100...and b=(4/5)b+a/2...a=a/2+c/3...three equations three variable solve for the answer

OpenStudy (anonymous):

Ok I I have those equations but I am not sure how to go from there ( I have never worked with three variables) @akash_809

OpenStudy (anonymous):

use c=100-a-b, put c in a=a/2+c/3 now u will have a equation in a nd b and u already have a equation in a and b that is b=(4/5)b+a/2...solve these two equations

OpenStudy (anonymous):

Okay thanks !

OpenStudy (anonymous):

np @edsgirl17